# From Fibers to Cells: Fourier-Based Registration Enables Virtual Cresyl Violet Staining From 3D Polarized Light Imaging

### v1

@@ Abstract
Comprehensive assessment of the various aspects of the brain’s microstructure requires the use of complementary imaging techniques. This includes measuring the spatial distribution of cell bodies (cytoarchitecture) and nerve fibers (myeloarchitecture). The gold standard for cytoarchitectonic analysis is light microscopic imaging of cell-body stained tissue sections. To reveal the 3D orientations of nerve fibers, 3D Polarized Light Imaging (3D-PLI) has been introduced as a reliable technique providing a resolution in the micrometer range while allowing processing of series of complete brain sections. 3D-PLI acquisition is label-free and allows subsequent staining of sections after 3D-PLI measurement. By post-staining for cell bodies, a direct link between fiber- and cytoarchitecture can potentially be established within one and the same section. However, inevitable distortions introduced during the staining process make a costly nonlinear and cross-modal registration necessary in order to study the detailed relationships between cells and fibers in the images. In addition, the complexity of processing histological sections for post-staining only allows for a limited number of samples. In this work, we take advantage of deep learning methods for image-to-image translation to generate a virtual staining of 3D-PLI that is spatially aligned at the cellular level. We use a supervised setting, building on a unique dataset of brain sections, to which Cresyl violet staining has been applied after 3D-PLI measurement. To ensure a high correspondence between both modalities, we address the misalignment of training data using Fourier-based registration methods. In this way, registration can be efficiently calculated during training for local image patches of target and predicted staining. We demonstrate that the proposed method enables prediction of a Cresyl violet staining from 3D-PLI, matching individual cell instances.
@@ 1 Introduction
To understand the organizational principles of the brain, complementary imaging techniques are used to highlight different aspects of brain architecture. Two important aspects of the microstructural organization are fiber- and cytoarchitecture Nieuwenhuys, (2013); Amunts and Zilles, (2015). While cytoarchitecture encompasses the spatial distribution and shape of cell bodies in the cerebral cortex and subcortical nuclei, fiber architecture refers to the course and composition of nerve fibers. However, cytoarchitecture and fiber-architecture are usually studied using different staining protocols, applied in different sections. Only a few protocols are available to combine cyto- and fiber staining in a single protocol, e.g., Luxol fast blue Klüver and Barrera, (1953), Bielschowsky Bielschowsky, (1904) or the triple staining by Novotny Novotny and and Novotny, (1977) and Nowotny (2009). While they allow visualizing cell bodies and fibers in one and the same section, they do not allow the tracing of axons and fiber bundles over long distances.
@@
3D-Polarized Light Imaging (3D-PLI) addresses this limitation. It is a microscopic imaging technique for evaluating the three-dimensional orientation of myelinated nerve fibers in entire, unstained histological brain sections Axer and Amunts, (2022); Axer et al., 2011a ; Axer et al., 2011b . The technique can achieve an in-plane resolution of 1.3 µm, capturing structures at the level of individual fibers and small fiber bundles. 3D-PLI has been used to gain insights into the architecture of nerve fibers in different brain regions, such as the human hippocampus Zeineh et al., (2017), the sagittal stratum Caspers et al., (2022) and the vervet monkey visual system Takemura et al., (2020). In addition, 3D-PLI has been used to validate fiber tractography algorithms and the interpretation of DW-MRI Caspers and Axer, (2019). 3D-PLI potentially allows joint imaging of fiber tracts and neuronal cell bodies Zeineh et al., (2017) due to scattering effects affecting light transmittance Menzel et al., (2020). However, this possibility has not yet been validated.
@@
Cytoarchitecture can be studied in histological sections of postmortem brains with Cresyl violet. The staining provides contrast due to staining of the rough endoplasmic reticulum. This allows to study cell shape, density and distribution, which vary between brain regions. Due to its high spatial resolution, microscopic analysis of histological sections is considered the gold standard to verify structural parcellations Amunts and Zilles, (2015). Recent advances in high-throughput scanning, data processing algorithms, and computational capacities have enabled the creation of 3D human brain atlases based on cytoarchitecture, such as BigBrain Amunts et al., (2013), the Allen Adult Human Brain Atlas (AAHA, Ding et al., 2016), and the Julich Brain probabilistic atlas Amunts et al., (2020).
@@
Since 3D-PLI relies solely on optical properties of the tissue, it is label-free and can be combined with a staining of the same tissue after its measurement. Post-staining, e.g. with Cresyl violet, enables a complementary visualization of neuronal cell bodies, potentially establishing a direct link between cytoarchitecture with 3D fiber-architecture. This requires, however, a complex histological processing, which limits the number of available samples, increases the risk of tissue damage and may lead to deformations of the section. To correct for deformation and artifacts in the images of the two modalities requires a nonlinear registration step. However, 3D-PLI and Cresyl violet stained tissue share only a few automatically identifiable cross-modal registration landmarks, such as blood vessels or morphological landmarks. I.e., post-staining of sections imaged with 3D-PLI is feasible, but technically challenging, and it hardly scales up.
@@
Therefore, we aim to train a deep neural network model to perform image-to-image translation from 3D-PLI to a Cresyl violet staining. Such an approach is often denoted as virtual histological staining, which refers to computational methods that generate color-coded images of biological tissue without the need for traditional staining techniques. The methods instead utilize optical properties of the tissue, such as birefringence, autofluorescence, scattering, or absorption to create images that emulate the appearance of stained tissue. A virtual Cresyl violet staining spatially aligned with 3D-PLI would allow an application of established tools for cytoarchitectonic analysis, such as automatic cell instance segmentation Upschulte et al., (2022), for 3D-PLI measurements. Furthermore, it would provide detailed registration landmarks for cross-modal registration Ounkomol et al., (2018), thereby offering the opportunity to perform joint acquisition of aligned fiber and cytoarchitecture at a larger scale than possible today.
@@
One of the earliest applications of label-free imaging for visual staining was Quantitative Phase Imaging (QPI, Curl et al., 2004). QPI measures the phase shift of light passing through a sample, producing high-resolution images that reveal the optical properties of the tissue. It was used to generate images of collagen fibers, red blood cells, and other tissue structures without staining Curl et al., (2004); Park et al., (2018). Later, machine-learning algorithms, especially generative models, have been trained to recognize and virtually stain different tissue structures in unstained images by performing image-to-image translation. They have successfully generated color-coded images of tissue that replicate the appearance of histological stainings, such as a virtual hematoxylin and eosin (H&E), Masson’s trichrome, and Jones’ stain from QPI of label-free tissue Rivenson et al., (2019), a transformation of H&E stained tissue into Masson’s trichrome, periodic acid-Schiff (PAS), or Jones’ stain de Haan et al., (2021); Yang et al., (2022). However, these stains are not very good at distinguishing the different components of the nervous tissue.
@@
Machine learning algorithms were also used to predict fluorescence-labeled images from transmitted-light z-stacks Ounkomol et al., (2018); Christiansen et al., (2018); Cross-Zamirski et al., (2022) or 3D fluorescence structures and a FluoroMyelin stain from bright-field and polarization images of brain slices Guo et al., (2020). The methods typically use cross-entropy Christiansen et al., (2018), mean absolute (L1) loss Guo et al., (2020), or style-related losses such as conditional generative adversarial network (GAN) loss Rivenson et al., (2019); de Haan et al., (2021); Cross-Zamirski et al., (2022). While including a GAN objective encourages prediction of realistic-looking images, it has no clear mechanism to preserve content when conditioned on a particular input image, and thus may introduce artificial structures Cohen et al., (2018). A combination with a pixel-wise reconstruction loss (e.g. L1 loss) mitigates this problem of GAN training and improves accuracy of predictions Isola et al., (2017).
@@
Since methods using paired training data for supervised image-to-image translation typically produce more accurate predictions than unpaired methods Zhu et al., (2017), a pixel-accurate alignment of training data is necessary. This requires virtual staining methods to either perform a costly registration step or directly acquire paired images. A paired acquisition with 3D-PLI, however, is not feasible and a lack of structural overlap, such as a sufficient number of visible cell instances between the investigated modalities, makes pixel-accurate registration challenging. Therefore, to alleviate the need for perfectly paired training data, we propose a supervised learning objective performing local online registration of training pairs combined with a translational-invariant style comparison. This allows us to train the model on imperfectly registered image pairs with strong content preservation as in paired image-to-image translation, while enabling realistic prediction of subtle structures like cell bodies (Fig. 1). The main contributions are as follows:
@@
  • We apply the matching of Gram matrix representations as a texture sensitive style loss for the virtual staining, as previously used for texture synthesis Gatys et al., (2015). Since the computation of Gram matrices is translation invariant, it allows a direct comparison of image statistics between coarsely aligned training examples. It therefore improves the accuracy of predicted cell instances over commonly used GAN style loss.
@@
  • An online registration head for improving registration accuracy of local image pairs after pre-alignment of larger tissue sections during training. We consider Fourier-based registration methods, which can be computed efficiently in real-time on modern GPU hardware.
@@ 2.1 Microscopic imaging of histological brain sections
We demonstrate the proposed virtual staining approach on a set of brain sections for which Cresyl violet staining has been performed after 3D-PLI measurement. The brain sample for this study was obtained from a healthy 2.4-year-old adult male vervet monkey (Wake Forest-ID 1818, Axer et al., 2020; Takemura et al., 2020; Menzel et al., 2022) in accordance with the Wake Forest Institutional Animal Care and Use Committee (IACUC #A11-219) and conforming the AVMA Guidelines for the Euthanasia of Animals. To obtain an undistorted volumetric reference, a T2 weighted MRI was acquired in-vivo one day prior to sacrifice. The brain was removed from the skull within 24 hours after flushing with phosphate-buffered saline and perfusion fixated with 4% buffered paraformaldehyde. It was stored for several weeks at -70°C in 20% glycerin solution for cryoprotection, and then sliced coronally with 60 µm section thickness using a large-scale cryostat microtome (Polycut CM 3500, Leica Microsystems, Germany). Blockface images of the frozen tissue block were taken with a CCD camera before cutting each brain section. The images were reconstructed into a 3D blockface volume to provide an undistorted reference for section realignment.
@@ 2.1.1 Image acquisition
Figure 2: Data modalities and registration challenges for training section 544. (A-C) 3D-PLI parameter maps: Transmittance $I_{T}$, retardation $\sin\delta$ and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$). Background pixels are masked for visualization purposes only. (D) Affine registered Cresyl violet staining. The pial surface of the 3D-PLI acquisition is shown as a contour plot in D for reference. Between both data acquisitions remains a nonlinear misalignment that cannot be resolved by a global affine transformation. At a local scale, the remaining misalignment is approximately linear. Yellow arrows indicate blood vessels that can be used as mutual registration landmarks for coarse alignment.
For 3D-PLI acquisition, brain sections were scanned using a polarizing microscope (LMP-1, Taorad, Germany) with 1.3 µm resolution Axer et al., 2011b ; Axer and Amunts, (2022). The focus level of the LMP-1 was manually adjusted to the center of the tissue for each section. Inside the LMP-1 microscope, sections were placed on a specimen stage between a rotating linear and a circular polarizer on top of an incoherent light source with a wavelength of 550 $\pm$ 5 nm. Images were taken by a CCD camera for nine equidistant rotation angles $\rho$ of the rotating linear polarizer, covering 180° of rotation. At each pixel, the measured intensity of the images followed a sinusoidal profile as
@@
Using harmonic Fourier analysis, parameter maps of transmittance ($I_{T}$), retardation ($\sin\delta$), and direction ($\varphi$) were obtained from the measurements, with a resolution of approximately 34,000 $\times$ 44,000 pixels per section, revealing their fine-grained nerve fiber architecture (Fig. 2A-C). After 3D-PLI acquisition, brain sections were washed, fixed and stained for cell bodies with Cresyl violet Nissl staining to reveal their cellular architecture. They were scanned using flatbed scanners at 1 µm in-plane resolution and saved as RGB color pictures with eight bit color depth, storing pixel values from 0 to 255 (Fig. 2D).
2.1.2 Tissue shrinkage correction

### v2

@@ Abstract
Comprehensive assessment of the various aspects of the brain’s microstructure requires the use of complementary imaging techniques. This includes measuring the spatial distribution of cell bodies (cytoarchitecture) and nerve fibers (myeloarchitecture). The gold standard for cytoarchitectonic analysis is light microscopic imaging of cell-body stained tissue sections. To reveal the 3D orientations of nerve fibers, 3D Polarized Light Imaging (3D-PLI) has been introduced as a reliable technique providing a resolution in the micrometer range while allowing processing of series of complete brain sections. 3D-PLI acquisition is label-free and allows subsequent staining of sections after 3D-PLI measurement. By post-staining for cell bodies, a direct link between fiber- and cytoarchitecture can potentially be established in the same section. However, inevitable distortions introduced during the staining process make a costly nonlinear and cross-modal registration necessary in order to study the detailed relationships between cells and fibers in the images. In addition, the complexity of processing histological sections for post-staining only allows for a limited number of such samples. In this work, we take advantage of deep learning methods for image-to-image translation to generate a virtual staining of 3D-PLI that is spatially aligned at the cellular level. We use a supervised setting, building on a unique dataset of brain sections, to which Cresyl violet staining has been applied after 3D-PLI measurement. To ensure high correspondence between both modalities, we address the misalignment of training data using Fourier-based registration. In this way, registration can be efficiently calculated during training for local image patches of target and predicted staining. We demonstrate that the proposed method can predict a Cresyl violet staining from 3D-PLI, resulting in a virtual staining that exhibits plausible patterns of cell organization in gray matter, with larger cell bodies being localized at their expected positions.
@@ 1 Introduction
To understand the organizational principles of the brain, complementary imaging techniques are used to highlight different aspects of brain architecture. Two important aspects of the microstructural organization are fiber- and cytoarchitecture nieuwenhuys2013; amunts2015. While cytoarchitecture encompasses the spatial distribution and shape of cell bodies in the cerebral cortex and subcortical nuclei, fiber architecture refers to the course and composition of nerve fibers. However, cytoarchitecture and fiber-architecture are usually studied using different staining protocols, applied in different sections. Only a few protocols are available to combine cyto- and fiber staining in a single protocol, e.g., Luxol fast blue kluver1953, Bielschowsky bielschowsky1904 or the triple staining by Novotny novotny1977. While they allow visualizing cell bodies and fibers in one and the same section, they lack information about 3D fiber orientations. As a result, they do not support the tracing of axons and fiber bundles over long distances.
@@
3D-Polarized Light Imaging (3D-PLI) addresses this limitation. It is a microscopic imaging technique for evaluating the three-dimensional orientation of myelinated nerve fibers in entire, unstained histological brain sections axer2022; axer2011; axer2011a. The technique can achieve an in-plane resolution of 1.3 µm, capturing structures at the level of individual fibers and small fiber bundles. 3D-PLI has been used to gain insights into the architecture of nerve fibers in different brain regions, such as the human hippocampus zeineh2017, the sagittal stratum caspers2022 and the vervet monkey visual system takemura2020. In addition, 3D-PLI has been used to validate fiber tractography algorithms and the interpretation of DW-MRI caspers2019. 3D-PLI potentially allows joint imaging of fiber tracts and neuronal cell bodies zeineh2017 due to diffraction patterns, differences in the density of birefringent material, and locally variable attenuation. However, this possibility has not yet been validated.
@@
Cytoarchitecture can be studied in histological sections of postmortem brains with Cresyl violet. The staining provides contrast due to staining of the rough endoplasmic reticulum. This allows to study cell shape, density and distribution, which vary between brain regions. Due to its high spatial resolution, microscopic analysis of histological sections is considered the gold standard to verify structural parcellations amunts2015. Recent advances in high-throughput scanning, data processing algorithms, and computational capacities have enabled the creation of 3D human brain atlases based on cytoarchitecture, such as BigBrain amunts2013, the Allen Adult Human Brain Atlas (AAHA, ding2016), and the Julich Brain probabilistic atlas amunts2020.
@@
Since 3D-PLI relies solely on optical properties of the tissue, it is label-free and can be combined with a staining of the same tissue after its measurement. Post-staining, e.g. with Cresyl violet, enables a complementary visualization of neuronal cell bodies, potentially establishing a direct link between cytoarchitecture with 3D fiber-architecture. This requires, however, a complex histological processing, which limits the number of available samples, increases the risk of tissue damage and may lead to deformations of the section. To correct for deformation and artifacts in the images of the two modalities requires a nonlinear registration step. However, 3D-PLI and Cresyl violet stained tissue share only a few automatically identifiable cross-modal registration landmarks, such as blood vessels or morphological landmarks. Therefore, post-staining of sections imaged with 3D-PLI is feasible but technically challenging. Since it does not scale efficiently to larger datasets, it is not applicable to whole-brain stacks involving thousands of sections.
@@
Therefore, we aim to train a deep neural network model to perform image-to-image translation from 3D-PLI to a Cresyl violet staining. Such an approach is often denoted as virtual histological staining, which refers to computational methods that generate color-coded images of biological tissue without the need for traditional staining techniques latonen2024. The methods instead utilize optical properties of the tissue, such as birefringence, autofluorescence, scattering, or absorption to create images that emulate the appearance of stained tissue. A virtual Cresyl violet staining spatially aligned with 3D-PLI would allow the use of established tools for cytoarchitectonic analysis, such as automatic cell instance segmentation upschulte2022, directly on 3D-PLI data. Furthermore, identification of cell bodies would provide detailed registration landmarks for cross-modal registration ounkomol2018, thereby offering the opportunity to perform joint acquisition of aligned fiber and cytoarchitecture at a larger scale than possible today. A virtual staining could, in principle, be applied to the whole brain. However, white matter regions remain challenging due to the predominance of glial cells, mainly oligodendrocytes forming the myelin sheath, which are not distinguishable from nerve fibers in the 3D-PLI signal. Therefore, in the present study, we focus our analysis on gray matter regions.
@@
One of the earliest applications of label-free imaging for visual staining was Quantitative Phase Imaging (QPI, curl2004). QPI measures the phase shift of light passing through a sample, producing high-resolution images that reveal the optical properties of the tissue. It was used to generate images of collagen fibers, red blood cells, and other tissue structures without staining curl2004; park2018. Later, machine-learning algorithms, especially generative models, have been trained to recognize and virtually stain different tissue structures in unstained images by performing image-to-image translation. They have successfully generated color-coded images of tissue that replicate the appearance of histological stainings, such as a virtual hematoxylin and eosin (H&E), Masson’s trichrome, and Jones’ stain from QPI of label-free tissue rivenson2019, a transformation of H&E stained tissue into Masson’s trichrome, periodic acid-Schiff (PAS), or Jones’ stain dehaan2021; yang2022. However, these stains are not very good at distinguishing the different components of the nervous tissue.
@@
Machine learning algorithms were also used to predict fluorescence-labeled images from transmitted-light z-stacks ounkomol2018; christiansen2018; cross-zamirski2022 or 3D fluorescence structures and a FluoroMyelin stain from bright-field and polarization images of brain slices guo2020. The methods typically use cross-entropy christiansen2018, mean absolute (L1) loss guo2020, or style-related losses such as conditional generative adversarial network (GAN) loss rivenson2019; dehaan2021; cross-zamirski2022. While including a GAN objective encourages prediction of realistic-looking images, it has no clear mechanism to preserve content when conditioned on a particular input image, and thus may introduce artificial structures cohen2018. A combination with a pixel-wise reconstruction loss (e.g. L1 loss) mitigates this problem of GAN training and improves accuracy of predictions isola2017.
@@
Since methods using paired training data for supervised image-to-image translation typically produce more accurate predictions than unpaired methods zhu2017; latonen2024, a pixel-accurate alignment of training data is desired. This requires virtual staining methods to either perform a costly registration step or directly acquire paired images. A paired acquisition with 3D-PLI, however, is not feasible and a lack of structural overlap, such as a sufficient number of visible cell instances between the investigated modalities, makes pixel-accurate registration challenging. Therefore, to alleviate the need for perfectly paired training data, we propose a supervised learning objective performing local online registration of training pairs combined with a translational-invariant style comparison. This allows us to train the model on imperfectly registered image pairs with strong content preservation as in paired image-to-image translation, while enabling realistic prediction of subtle structures like cell bodies (Fig.˜1).
The main contributions of our method are the following:
@@
  • We apply the matching of Gram matrix representations as a texture sensitive style loss for the virtual staining, as previously used for texture synthesis gatys2015a. Since the computation of Gram matrices is translation invariant, it allows a direct comparison of image statistics between coarsely aligned training examples. It therefore improves the accuracy of predicted cell instances over commonly used GAN style loss.
@@
  • An online registration head for improving registration accuracy of local image pairs after pre-alignment of larger tissue tiles during training. We consider Fourier-based registration methods, which can be computed efficiently in real-time on modern GPU hardware.
@@ 2.1 Microscopic imaging of histological brain sections
We demonstrate the proposed virtual staining approach on a set of brain sections for which Cresyl violet staining has been performed after 3D-PLI measurement. The brain sample for this study was obtained from a healthy 2.4-year-old adult male vervet monkey (Wake Forest-ID 1818, axer2020vervet; takemura2020) in accordance with the Wake Forest Institutional Animal Care and Use Committee (IACUC #A11-219) and conforming the AVMA Guidelines for the Euthanasia of Animals. To obtain an undistorted volumetric reference, a T2 weighted MRI was acquired in-vivo one day prior to sacrifice. The brain was removed from the skull within 24 hours after flushing with phosphate-buffered saline and perfusion fixated with 4% buffered paraformaldehyde. It was stored for several weeks at -70°C in 20% glycerin solution for cryoprotection, and then sliced coronally with 60 µm section thickness using a large-scale cryostat microtome (Polycut CM 3500, Leica Microsystems, Germany). Blockface images of the frozen tissue block were taken with a CCD camera before cutting each brain section. The images were reconstructed into a 3D blockface volume to provide an undistorted reference for section realignment schober2015.
@@ 2.1.1 Image acquisition
Figure 2: Data modalities and registration challenges for training section 544. (A-C) 3D-PLI parameter maps: Transmittance, retardation and fiber orientation in HSV color space (hue: fiber direction; saturation/brightness: retardation). Background pixels are masked for visualization purposes only. (D) Affine registered Cresyl violet staining. The pial surface of the 3D-PLI acquisition is shown as a contour plot in D for reference. Between both data acquisitions remains a nonlinear misalignment that cannot be resolved by a global affine transformation. At a local scale, the remaining misalignment is approximately linear. Yellow arrows indicate blood vessels that can be used as mutual registration landmarks for coarse alignment.
For 3D-PLI acquisition, brain sections were scanned using a polarizing microscope (LMP-1, Taorad, Germany) with 1.3 µm resolution axer2011a; axer2022. The focus level of the LMP-1 was manually adjusted to the center of the tissue for each section. Inside the LMP-1 microscope, sections were placed on a specimen stage between a rotating linear and a circular polarizer on top of an incoherent light source with a wavelength of 550 $\pm$ 5 nm. Images were taken by a CCD camera for nine equidistant rotation angles $\rho$ of the rotating linear polarizer, covering 180° of rotation. At each pixel, the measured intensity of the images followed a sinusoidal profile as
@@
Using harmonic Fourier analysis, parameter maps of transmittance ($I_{T}$), retardation ($\sin\delta$), and fiber direction ($\varphi$) were obtained from the measurements, with an image size of approximately 34,000 $\times$ 44,000 pixels per section, revealing their fine-grained nerve fiber architecture (Fig.˜2A-C). Each parameter map was stored in a separate HDF5 file as uncompressed 32-bit floating-point single-channel image.
After 3D-PLI acquisition, brain sections were washed, fixed and stained for cell bodies with Cresyl violet Nissl staining to reveal their cellular architecture. Whole-slide flat scans (single-plane) were performed using a Huron TissueScope LE120 high-throughput scanner at 1 µm in-plane resolution (Fig.˜2D). The resulting images were saved as RGB color images with eight bit color depth (pixel values ranging from 0 to 255) in uncompressed BigTIFF format.
2.1.2 Optical effects of cell bodies on the 3D-PLI signal
While 3D-PLI was primarily developed to map fiber orientations, cell bodies contribute to the measured signal as well. In the following, we summarize how absorption, diffraction, birefringence and scattering effects of cells are represented in 3D-PLI parameter maps.
Previous work reported that larger cell bodies appear as dark spots in transmittance maps zeineh2017, which encode light extinction caused by any material along the optical path. However, in the present transmittance maps, cell bodies are not distinguished by higher absorption relative to the surrounding fiber architecture (Fig.˜2A). This is likely because their membranes remained intact due to the short postmortem time before tissue fixation.
Diffraction significantly impacts the 3D-PLI signal at the given wavelength and resolution. In cortical regions, diffraction can cause pixels of transmittance maps inside the tissue to appear brighter than the background, particularly along sharp edges such as the walls of cell bodies and blood vessels. The intensity of these diffraction patterns on the transmittance map depends on the level of the focal plane of the objective lens.
Another relevant effect is observed in retardation maps, which encode the average amount and orientation of birefringent material within each tissue voxel, primarily collagen and myelinated nerve fibers. Since cell bodies contain significantly less birefringent material than surrounding fibers, their presence causes local attenuation of the retardation appearing as dark patches between cortical fibers (Fig.˜2B).
Light scattering is ubiquitous in polarized light bright-field transmission microscopy, but can generally be treated as random background noise. It does not appear to systematically affect either transmittance or retardation, with the only known exception being increased scattering at steep fibers running approximately perpendicularly to the section plane and darkening the transmittance significantly.
2.1.3 Tissue shrinkage estimation
Deformations induced by histological processing include shrinkage or swelling of brain tissue. To correct these effects, the extent of shrinkage can be estimated from the ratio between the histologically processed and true brain volume, which can be represented by the fresh weight of the whole brain with an estimated mean specific density amunts2005; amunts2013 or an MRI reference wagstyl2020. In this work, we perform 2D segmentation of cell bodies in 3D-PLI parameter maps and measure their in-plane areas. To ensure comparability of cell sizes across studies, we estimate 2D shrinkage factors using a postmortem MRI of the same brain as a reference.